# Longest cycles in vertex-transitive and highly connected graphs

### v1

@@ 3. Reduction to computer search
By Menger’s Theorem, there are at least $\textup{ex}(2k,\{Q_{3}^{+},K_{3,3}\})+1$ vertex-disjoint paths with one endpoint in $V(C_{1})\setminus V(C_{2})$ and one endpoint in $V(C_{2})\setminus V(C_{2})$. Fix such a collection $\mathcal{P}$. We also fix total orderings on $C_{1}$ and $C_{2}$ by choosing a smallest element and a direction.
@@
Since $|E|=|\mathcal{P}|\geq\textup{ex}(2k,\{Q_{3}^{+},K_{3,3}\})+1$, it follows that the bipartite graph $H=(X,Y,E)$ contains either $Q_{3}^{+}$ or $K_{3,3}$. Thus, by Lemma 3.4, $(\mathcal{G}_{\tau},W)$ contains a long good cycle $C^{\tau}$. The vertices $v^{\tau}_{1},\dots,v^{\tau}_{\ell}$ of $C^{\tau}$ correspond to vertices $v_{1},\dots,v_{\ell}$ of $G$. The path which was used to define the edge $e=v_{i}^{\tau}v_{i+1}^{\tau}$ of $C^{\tau}$ has length $w(e)$ and is between $v_{i}$ and $v_{i+1}$. Therefore, we obtain a closed walk $C$ in $G$ of length $w(C^{\tau})$. Since $C^{\tau}$ is long, $w(C^{\tau})>w(C_{1}^{\tau})$ and the length of $C$ is strictly larger than the length of $C_{1}$. What remains to show is that $C$ is in fact a cycle. The vertices $v_{1},\dots,v_{\ell}$ are distinct because the vertices $v^{\tau}_{1},\dots,v^{\tau}_{\ell}$ are distinct, but we need to show the paths between them do not intersect internally. The paths corresponding to cross edges are chosen to be internally vertex-disjoint and only intersect $V(C_{1})\cup V(C_{2})$ in their endpoints. For $i\in\{1,2\}$, the paths corresponding to edges from $C_{i}^{\tau}$ correspond to internally disjoint arcs of $C_{i}$. It remains to rule out intersections between the arcs of $C$ on $C_{1}$ and $C_{2}$: By the definition of good, there exists some $i\in\{1,2\}$ such that all arcs of $C_{i}$ that are used by $C$ are part of a connected component of $C_{i}-V(C_{3-i})$ and thus do not intersect any of the arcs that $C$ uses from $C_{3-i}$. This shows $C$ is indeed a cycle, contradicting $C_{1}$ being the longest cycle.
@@ 5. Further remarks
One direction towards improving the bounds in Lemma 1.4 is to prove a result analogous to Lemma 3.4 for a graph $H$ with a better extremal number than $Q_{3}^{+}$. In fact, Chen, Faudree and Gould [12] already took a similar approach back in 1998 and proved a result analogous to Lemma 3.4 for $H=K_{3,257}$ without the aid of a computer, via a result analogous to our Lemma 3.1, Erdős-Szekeres and the pigeonhole principle. By computer search, we found that Lemma 3.4 is false when $Q_{3}^{+}$ is replaced by $Q_{3}$: there are (weighted) configuration graphs on an ordered bipartite graph of $Q_{3}$ without long good cycles. Indeed, these can be obtained with only slight modifications to the code in [25], see Figure 3 for an example.
21212121111131313131333333331111707070701111333371717171535353531313131338383838292929293939393981818181444428282828333333331111111135353535111115151515696969691111111111111111111111111111111111111111121212121111
Figure 3. Example of a weighted configuration graph on an ordered bipartite graph with $H=Q_{3}$, for which there does not exist any long good cycle. The weight edges are specified as edge labels. Dangerous edges are dashed.
@@
As mentioned previously in this paper, one can observe that $G$ is in fact $2$-connected and that every two longest paths and every two longest cycles in $G$ must meet. A key observation is that if a mapping $\phi:V(G)\rightarrow V(G)$ is drawn uniformly at random from the set of all automorphisms of $G$, then, since $G$ is vertex-transitive, for every fixed vertex $u\in V(G)$, its image $\phi(u)$ is uniformly distributed among all vertices of $G$.
@@
$=\sum_{v\in V(G)}\prod_{i=1}^{k}{\sum_{u\in V(P)}\mathbb{P}[\phi_{i}(u)=v]}=n\cdot\left(\frac{|V(P)|}{n}\right)^{k}=\frac{|V(P)|^{k}}{n^{k-1}}.$
@@
A famous question asked by Gallai [23] in 1966 is whether $\mathrm{lpt}(G)=1$ for every connected graph $G$, i.e., whether all longest paths meet in one vertex. This was answered in the negative by Walther [43], who constructed a connected graph $G$ on $25$ vertices with $\mathrm{lpt}(G)=2$. A smaller counterexample on $12$ vertices was found independently by Walther and Zamfirescu [44, 48]. Later, Grünbaum [27] constructed a connected graph $G$ on $324$ vertices with $\mathrm{lpt}(G)=3$, and amazingly, this remains the best known lower bound for connected graphs to this date. In particular, it is still an open question, raised by both Walther and Zamfirescu [47], whether there exists an absolute constant $c>0$ such that $\mathrm{lpt}(G)\leq c$ for every connected graph $G$. Note that this would immediately yield a lower bound of $\Omega(n)$ on the maximum length of paths in vertex-transitive graphs, but it remains wide open. In fact, all the known general upper bounds for longest path and cycle transversals in connected and $2$-connected graphs that are known thus far grow polynomially in $n$, see the recent sequence of papers [34, 37, 39] on this problem. The current state of the art bounds are $\mathrm{lpt}(G)\leq 5n^{2/3}$ for all connected $n$-vertex graphs and $\mathrm{lct}(G)\leq 5n^{2/3}$ for all $2$-connected $n$-vertex graphs by Kierstead and Ren [34]. Unfortunately, these only provide lower bounds of order $\Omega(n^{1/3})$ on longest paths and cycles in vertex-transitive graphs, but improving the upper bounds on $\mathrm{lpt}(G)$ and $\mathrm{lct}(G)$ to $n^{o(1)}$, for instance, would be very interesting.

### v2

@@ 99 $\{a_{4},b_{4}\}$ $2052$ $2052$
By Menger’s Theorem, there are at least $\textup{ex}(2k,\{Q_{3}^{+},K_{3,3}\})+1$ vertex-disjoint paths with one endpoint in $A$ and one endpoint in $B$. Fix such a collection $\mathcal{P}$. We also fix total orderings on $C_{1}$ and $C_{2}$ by choosing a smallest element and a direction.
@@
Since $|E|=|\mathcal{P}|\geq\textup{ex}(2k,\{Q_{3}^{+},K_{3,3}\})+1$, it follows that the bipartite graph $H=(X,Y,E)$ contains either $Q_{3}^{+}$ or $K_{3,3}$. Thus, by Lemma 3.4, $(\mathcal{G}_{\tau},W)$ contains a good long cycle $C^{\tau}$. The vertices $v^{\tau}_{1},\dots,v^{\tau}_{\ell}$ of $C^{\tau}$ correspond to vertices $v_{1},\dots,v_{\ell}$ of $G$. The path which was used to define the edge $e=v_{i}^{\tau}v_{i+1}^{\tau}$ of $C^{\tau}$ has length $w(e)$ and is between $v_{i}$ and $v_{i+1}$. Therefore, we obtain a closed walk $C$ in $G$ of length $w(C^{\tau})$. Since $C^{\tau}$ is long, $w(C^{\tau})>w(C_{1}^{\tau})$ and the length of $C$ is strictly larger than the length of $C_{1}$. What remains to show is that $C$ is in fact a cycle. The vertices $v_{1},\dots,v_{\ell}$ are distinct because the vertices $v^{\tau}_{1},\dots,v^{\tau}_{\ell}$ are distinct, but we need to show the paths between them do not intersect internally. The paths corresponding to cross edges are chosen to be internally vertex-disjoint and only intersect $V(C_{1})\cup V(C_{2})$ in their endpoints. For $i\in\{1,2\}$, the paths corresponding to edges from $C_{i}^{\tau}$ correspond to internally disjoint arcs of $C_{i}$. It remains to rule out intersections between the arcs of $C$ on $C_{1}$ and $C_{2}$: By the definition of good, there exists some $i\in\{1,2\}$ such that all arcs of $C_{i}$ that are used by $C$ are part of a connected component of $C_{i}-V(C_{3-i})$ and thus do not intersect any of the arcs that $C$ uses from $C_{3-i}$. This shows $C$ is indeed a cycle, contradicting $C_{1}$ being the longest cycle.
@@ 5. Further remarks
One direction towards improving the bounds in Lemma 1.4 is to prove a result analogous to Lemma 3.4 for a graph $H$ with a better extremal number than $Q_{3}^{+}$. In fact, Chen, Faudree and Gould [14] already took a similar approach back in 1998 and proved a result analogous to Lemma 3.4 for $H=K_{3,257}$ without the aid of a computer, via a result analogous to our Lemma 3.1, the Erdős-Szekeres theorem and the pigeonhole principle. The perhaps most natural attempt is to consider the complete bipartite graphs $K_{2,t}$, since their extremal functions satisfy $\textup{ex}(n,K_{2,t})=\Theta(n^{3/2})\ll n^{8/5}$. However, one can show that for every fixed integer $t\in\mathbb{N}$ there exist (weighted) configuration graphs on an ordered bipartite graph of $K_{2,t}$ without good long cycles. Similarly, by computer search, we found that Lemma 3.4 is false when $Q_{3}^{+}$ is replaced by $Q_{3}$: there are (weighted) configuration graphs on an ordered bipartite graph of $Q_{3}$ without good long cycles. Indeed, these can be obtained with only slight modifications to the code in [27], see Figure 3 for an example.
2121113131333311707011337171535313133838292939398181442828333311113535111515696911111111111111111111121211
Figure 3. Example of a weighted configuration graph on an ordered bipartite graph with $H=Q_{3}$, for which there does not exist any good long cycle. The weight edges are specified as edge labels. Dangerous edges are dashed.
@@
It is easy to observe that $G$ must in fact be $2$-connected: Removing a leaf of a spanning tree of $G$ leaves a connected graph, and since $G$ is vertex-transitive, this means that remvoving any vertex of $G$ leaves a connected graph. Hence, by Observation 2.1 any two longest paths and any two longest cycles in $G$ intersect each other. A key observation is that if a mapping $\phi:V(G)\rightarrow V(G)$ is drawn uniformly at random from the set of all automorphisms of $G$, then, since $G$ is vertex-transitive, for every fixed vertex $u\in V(G)$, its image $\phi(u)$ is uniformly distributed among all vertices of $G$.
@@
$\sum_{v\in V(G)}\prod_{i=1}^{k}{\sum_{u\in V(P)}\mathbb{P}[\phi_{i}(u)=v]}=n\cdot\left(\frac{|V(P)|}{n}\right)^{k}=\frac{|V(P)|^{k}}{n^{k-1}}.$
@@
A famous question asked by Gallai [25] in 1966 is whether $\mathrm{lpt}(G)=1$ for every connected graph $G$, i.e., whether all longest paths meet in one vertex. This was answered in the negative by Walther [45], who constructed a connected graph $G$ on $25$ vertices with $\mathrm{lpt}(G)=2$. A smaller counterexample on $12$ vertices was found independently by Walther and Zamfirescu [46, 50]. Later, Grünbaum [29] constructed a connected graph $G$ on $324$ vertices with $\mathrm{lpt}(G)=3$, and amazingly, this remains the best known lower bound for connected graphs to this date. In particular, it is still an open question, raised by both Walther and Zamfirescu [49], whether there exists an absolute constant $c>0$ such that $\mathrm{lpt}(G)\leq c$ for every connected graph $G$. Note that this would immediately yield a lower bound of $\Omega(n)$ on the maximum length of paths in vertex-transitive graphs. The question of Walther and Zamfirescu is currently out of reach, and in fact, all the known general upper bounds for longest path and cycle transversals in connected and $2$-connected graphs grow polynomially in $n$, see the recent sequence of papers [36, 39, 41] on this problem. The current state of the art bounds are $\mathrm{lpt}(G)\leq 5n^{2/3}$ for all connected $n$-vertex graphs and $\mathrm{lct}(G)\leq 5n^{2/3}$ for all $2$-connected $n$-vertex graphs by Kierstead and Ren [36]. Unfortunately, these only provide lower bounds of order $\Omega(n^{1/3})$ on longest paths and cycles in vertex-transitive graphs, but improving the upper bounds on $\mathrm{lpt}(G)$ and $\mathrm{lct}(G)$ to $n^{o(1)}$, for instance, would be very interesting.